OpenStudy (anonymous):

Find A and B if x+16=A(x+2)+B(2x-3) is an identity.

hartnn (hartnn):

just replace x by 3/2 to get the value of A replace x by -2 to get the value of B what you get ?

hartnn (hartnn):

Note when you put x=3/2, the term multiplied with B becomes 0 thats the reason you could find the value of A directly.

OpenStudy (anonymous):

where did you get the 3/2?

OpenStudy (anonymous):

yey! A=5 B=-2? right?

hartnn (hartnn):

yes, thats correct :) i got 3/2 from 2x-3 =0 ----> 2x=3---->x=3/2 because , to find A, i wanted the term with B=0, so i equated 2x-3 to 0, which eliminated B.

hartnn (hartnn):

did you get that ^ ?

OpenStudy (anonymous):

x+16=A(x+2)+B(2x-3) x+16=Ax+2Bx+2A-3B x+16=(A+2B)x+(2A-3B) Therefore, (A+2B)x=x; 2A-3B=16 That is: A+2B=1 2A-3B=16 so we got A=5 B=-2

OpenStudy (anonymous):

@hartnn Thank you! :)